Thorium-234 has a half-life of 24.1 days. If a sample of thorium 234 has a mass of 6.4 x 10 -12 g, how
much of the sample is left after 72.3 days
1
Expert's answer
2020-11-20T10:05:36-0500

No=6.4x10^-12

T1/2=24.1

T=72.3

N=?

N=No(1/2)^t/t1/2= 6.4x10^-12 x (1/2)^ 72.3/24.1 = 8x10^-13.